OpenStudy (pottersheep):

Why can't -2 = x^(-1/2) ? PLEASE HELP!

OpenStudy (saifoo.khan):

\[\Huge x^{-\frac 12}=\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

x^(-1/2)=1/(x^(1/2))

OpenStudy (pottersheep):

uh huh

OpenStudy (pottersheep):

that still doesnt answer my quetion tho :s

OpenStudy (anonymous):

\[\sqrt{x}\] is positive, 1/positive is positive also.